# Improving Entropy-Based Test-Time Adaptation from a Clustering View

### v5

@@ Abstract
Domain shift is a common problem in the realistic world, where training data and test data follow different data distributions. To deal with this problem, fully test-time adaptation (TTA) leverages the unlabeled data encountered during test time to adapt the model. In particular, entropy-based TTA (EBTTA) methods, which minimize the prediction’s entropy on test samples, have shown great success. In this paper, we introduce a new perspective on the EBTTA, which interprets these methods from a view of clustering. It is an iterative algorithm: 1) in the assignment step, the forward process of the EBTTA models is the assignment of labels for these test samples, and 2) in the updating step, the backward process is the update of the model via the assigned samples. Based on the interpretation, we can gain a deeper understanding of EBTTA. Accordingly, we offer an alternative explanation for why existing EBTTA methods are sensitive to initial assignments, nearest neighbor information, outliers, and batch size. This observation can guide us to put forward the improvement of EBTTA. We propose to use robust label assignment, locality-preserving constraint, sample selection, and gradient accumulation to alleviate the above problems. Experimental results demonstrate that our method can achieve consistent improvements on various datasets. Code is provided in the supplementary material.
@@ Introduction
Despite the success of EBTTA, the performance of EBTTA would still vary greatly across different situations, e.g., Zhao et al. (Zhao et al. 2023) had shown that TENT holds empirical sensitivity and batch dependency through extensive experiments. Hence, it is necessary to gain a more thorough understanding of EBTTA, and then come up with a better solution.
In this paper, we provide a new perspective to better understand EBTTA through a clustering view. Clustering (Guo et al. 2017) is a traditional and effective way of unsupervised learning. The entropy-based methods and clustering both belong to unsupervised learning, where entropy minimization tries to reduce uncertainty about which category the sample belongs to, and clustering aims to map data into different clusters. Observing that, we show that entropy minimization and clustering have a lot in common, which makes it possible to use a clustering view to revisit the EBTTA. That is we can view the forward process in entropy minimization as the assignment step in clustering and the backward process (loss back-propagation and gradient descent) as the center updating step in clustering (see Section Interpretation for Entropy-Based Methods for detailed discussions).
Based on the interpretation, we can set the stage for improving the EBTTA. We first show that the entropy loss reduces the uncertainty by simply increasing the probability of the class with the largest value for each sample. According to this observation, several questions can be easily answered when we use the clustering view to revisit EBTTA, e.g., the initial assignments, nearest neighbor information, outliers, and batch dependency problems. With the analysis, we propose robust label assignment, locality-preserving constraint, sample selection, and gradient accumulation three simple modules to further improve the EBTTA. In summary, we propose to interpret the EBTTA from the view of clustering, which can guide us to develop better approaches for EBTTA.
@@
  • We revisit EBTTA from a clustering view: the forward process can be interpreted as the label assignment while the backward process of EBTTA can be interpreted as the update of centers in clustering methods.
  • The clustering view can guide us to put forward the improvement of EBTTA, where the robust label assignment, locality-preserving constraint, sample selection, and gradient accumulation are proposed to improve EBTTA.
  • We evaluate our method on various benchmark datasets, which demonstrates our method can gain consistent improvements over existing EBTTA.
@@
Fully test-time adaptation (TTA) aims to solve the problem of domain shift, where the training (source) domain and the test (target) domain follow different distributions. In order to fulfill this target, TTA online adapts the model using test samples, and a large amount of TTA methods have been proposed. For example, T3A (Iwasawa and Matsuo 2021) and Wang et al. (Wang et al. 2023) proposed to compute prototypes for each class and online updates the prototypes during adaptations; TeSLA (Tomar et al. 2023), RoTTA (Yuan, Xie, and Li 2023) and RMT (Döbler, Marsden, and Yang 2023) enforced the consistency between the teacher and student model; TIPI (Nguyen et al. 2023) and MEMO (Zhang, Levine, and Finn 2022) proposed to learn a model invariant of transformations.
The entropy-based test-time adaptation is a simple yet effective practice in the test-time adaptation, which minimizes the prediction’s entropy of test samples to adapt the model at test time. The most representative work of EBTTA is TENT (Wang et al. 2021), which conducted entropy minimizing via stochastic gradient descent and only modulated the BN layers. Then, several methods have been proposed to improve TENT. For example, ETA (Niu et al. 2022) and SAR (Niu et al. 2023) also minimized prections’s entropy but only selects the reliable samples as inputs. SHOT (Liang, Hu, and Feng 2020) augmented TENT by adding a divergence loss. AdaContrast (Chen et al. 2022) uses nearest neighbor information to generate more correct pseudo labels. CoTTA (Wang et al. 2022) used TENT to implement continual TTA, where the target domain changes sequentially. Despite the success of EBTTA, several works have found that the performance of EBTTA may sharply drop in some cases, e.g., when adapting on high-entropy samples (Niu et al. 2022) or when batch size is small (Wang et al. 2021; Nguyen et al. 2023). To explain this, we propose to view EBTTA as a clustering method and further make improvements on EBTTA based on the clustering view.
@@
Deep Spectral Clutsering
@@
The classical spectral clustering (Ng, Jordan, and Weiss 2001) stated above needs the calculation for eigenvectors thus lacks scalability. Deep spectral clustering (Shaham et al. 2018) is proposed to tackle this problem by learning the spectral embedding with a deep model using the locality-preserving constraint (Huang et al. 2014) as loss function. This approximation to classical spectral clustering can get better scalability and generalization ability with the cost of accuracy (Shaham et al. 2018).
@@
Mini-Batch K-Means: We give a brief description of mini-batch k-means (Bottou and Bengio 1994; Béjar Alonso 2013), which is used to interpret EBTTA methods. In the mini-batch k-means, the centers are first initialized, and then the A (assign) and U (update) steps are conducted alternately .
@@
U-step: With the probability $P(x)$ for each sample, EBTTA uses the entropy loss to update the model. This process is similar to the clustering methods that update the centers. In the clustering methods, the $i$-th center is updated by averaging features of the samples that were assigned to the $i$-th center in the A-step. We first give a lemma, which shows that the entropy loss also aims to further increase the samples’ probability belonging to the $i$-th class, where these samples have the largest probability of the $i$-th class in the forward process. The backward process of EBTTA is very similar to the U-step of clustering, that is both methods use the assigned samples to update the model/centers.
@@
The entropy loss would increase the probability of the class with the largest value, and decrease the sum of probabilities of the other classes.
@@
Table 1: TENT collapses when batch size is 1.
Remark 1: As shown in Table  1, TENT collapses when batch size is 1 and is worse than SOURCE method which only uses source model without adaptation. Using Lemma 1 can explain why EBTTA collapses when batch size is 1 and why EBTTA successes when batch size is larger than 1. Here follows the explanation.
@@
Lemma 1 means that if we get only one sample for each iteration (batch size = 1), then the model only tends to make the largest output probability predicted by source model larger until the output probability becomes one-hot as Figure 1 shows. This procedure in fact does not change the classification result. Therefore, when the batch size is 1, using TENT is ineffective as Table 1 shows.
However, if batch size is large than 1, then the gradients of multiple samples from target domain will be averaged to update a shared model. Hence, the largest probability of the data point may be changed due to the averaged gradients. This is similar to the clustering method. For example, $x_{1}$ is assigned to the positive class in the current iteration. After the update of cluster centers, $x_{1}$ may possibly be assigned to the negative class in the next iteration.
From the view of clustering, we can gain a more thorough understanding of EBTTA. Several questions can be more easily answered.
@@
Q1: What information does the source model provide during adaptation?
@@
In the TTA method, the source model is used as an initial model to be updated. From the clustering view, the source model also provides the initial probabilities belonging to the $K$ classes. As we had discussed above, the entropy loss tries to further increase the largest probability of each test sample. Thus the initial source model is very important, and we can achieve excellent performance with good initial probabilities, otherwise the entropy loss may hurt the performances. This is similar to clustering methods that are also sensitive to the initial cluster centers.
Q2: Why do some TTA works use the nearest neighbor information?
@@
Several TTA works like AdaContrast (Chen et al. 2022) and TAST (Jang, Chung, and Chung 2023) use nearest neighbor information to improve the adaptation ability. These works are under the assumption that a sample and its nearest neighbor in the feature space are likely to share the same label under the domain shift. The effectiveness of these works can be easily explained from the view of spectral clustering which also considers nearest neighbor information: using spectral clustering can make the clusters robust to complex scenarios (Liu and Han 2018). For example, if data points in a cluster after domain shift follows a non-Gaussian distribution, then it would a good choice to use spectral clustering to partition clusters.
Q3: Why do some EBTTA works need to select low-entropy samples?
@@
Several methods, e.g., ETA (Niu et al. 2022) and SAR (Niu et al. 2023), had proposed to use only the low-entropy samples to update the model, and the high-entropy samples were removed. The effectiveness of this technique has been verified by experiments but a detailed analysis is missing. From the clustering view, we can easily give an explanation: outliers should be suppressed. In Remark 1, we had shown that the gradients will be averaged to update the model. When the directions of the gradients are wrong (caused by the outliers), they may cause more mistakes in the updated model.
Q4: Why is batch size so important for EBTTA?
@@
The EBTTA methods are sensitive to the batch size (Wang et al. 2021; Nguyen et al. 2023). This can also be easily answered using the clustering view. As the deep model is updated via the averaged assigned samples’ gradients, more samples can provide more stable gradients. That is, when the batch size is small, the variance will be large. This is similar to the clustering methods, where the centers are updated by averaging features of assigned samples. When the number of samples is small, the calculated centers also will have large variance.
@@
# Locality Preserving Constraint
@@
Test-Time Clustering (TTC)
Based on the observations that EBTTA is sensitive to initial assignments, nearest neighbor information, outliers, and batch size, we propose to improve TENT from these four aspects accordingly. First, we suggest to obtain a more robust initial label assignment in order to get high-quality initial probabilities. Second, we propose to add the locality-preserving constraint to approximate spectral clustering. Third, we propose to select test samples to alleviate the negative effects of outliers. Finally, we recommend to use gradient accumulation to overcome the problem of small batch size.
Robust Label Assignment (RLA)
As discussed in Q1, we can see that label assignment in the A-step is very important for the EBTTA, which also provides a good initial probability for each test sample.
$\bar{f}(x)=\frac{f(x)+f(aug(x))}{2}$ (7)
$aug(x)=horizontal\_flip(x).$ (8)
This is because a weakly-augmented unlabeled sample can preserve more information about class label than the strongly augmented functions (Yang et al. 2022). Note that stop-gradient operation is applied to $f(aug(x))$ in Eq. (7), which can save computational resources.
Now the entropy loss for a single sample can be calculated as:
Locality-Preserving Constraint (LPC)
Spectral clustering is an improvement of k-means, which considers the similarity between samples in the same class. It uses the spectral embedding as data points for cluster partitioning. We add the locality-preserving constraint to generate spectral embedding (Huang et al. 2014; Yang et al. 2019; Shaham et al. 2018). Specifically, the constraint loss for a single sample $x$ can be formulated as:
Sample Selection (SS)
$H_{m}=\frac{1}{N}\sum_{x\in X}H(x).$ (11)
The selecting function $\mathcal{L}(x)$ can be formulated as:
Gradient Accumulation (GA)
The pseudo-code of our method is shown in Algorithm 1.
@@
The comparison results are shown in Table 2, Table 3 and Table 4. TENT w/ GA means TENT using the gradient accumulation. The results show that our method can achieve consistent improvements w.r.t. accuracy on all three datasets. The average (avg) accuracy is calculated on all corruption types. More detailed analysis is as follows.
CIFAR-10-C. The results are shown in Table 2. We can see that TTC achieves the best average accuracy 82.3%, which is 0.8% superior to the second-best method TENT. Besides, TTC performs better than TENT on most corruption types. Specifically, TTC is 2.3% superior to TENT w.r.t. impulse noise, which is a non-Gaussian transformation. Since TENT is the most relevant work, the comparison results indicate that the proposed improvements are effective.
CIFAR-100-C. As shown in Table 3, TTC is 1.9% better than TENT. Besides, TENT w/ GA performs the best, which is 0.7% better than NORM. It’s noting that NORM has achieved a good result, but using gradient accumulation can make TENT a better one.
ImageNet-C. The results of Table 4 show that TTC achieves the best average accuracy 42.8%, and the accuracy of the second-best model TENT is 41.8%. Note that previous entropy-based methods achieve excellent performance, even so, our method performs better than them.
@@
We conduct experiments to analyze the improvements caused by different components of our method individually. Specifically, we use TENT as the base model. We apply different components individually to TENT and the average accuracies for datasets are shown in Table 7. We shall see that Locality-Preserving Constraint (LPC) can consistently improve the performance on all datasets. Robust Label Assignment (RLA) can bring about improvements for TENT on CIFAR-10-C and CIAFR-100-C. Sample Selection (SS) can bring about improvements for TENT on both CIFAR-10-C and ImageNet-C. Gradient accumulation (GA) can make significant improvements on CIFAR-100-C. Please note that the batch size in Table 7 is 100, which is not too small. Thus it may be the reason why TENT+GA has similar result to TENT on CIFAR-10-C and ImageNet-C dataset. Detailed analysis of GA for different batch sizes can be found in Section Analysis of Different Batch Sizes, where GA can boost the performance significantly when the batch size is small.
@@
We conduct experiments to explore the effect of different augmentation functions $aug(x)$ in Eq.(7). The augmentation functions are horizontal flip (HorFlip), vertical flip (VerFlip), random rotation (RandRot), random resized crop (RRCrop), and Gaussian noise (GausNoise). From the average accuracies for datasets in Table 6, we shall see that horizontal flip provides a better pseudo-label and makes the performance better than other augmentation functions applied to TTC.
@@
We further conduct an ablation study to explore the effect of the batch size, and the average accuracies on datasets are shown in Figure 3. We shall see that all four methods perform better when batch size increases. Moreover, when applying gradient accumulation (GA) for small batch sizes, we can get a large performance improvement for TENT and TTC. For example, when the batch size is 2, TENT only gets 17.1% accuracy on CIFAR-10-C. However, when applying GA, we can get an improvement of 43.6% on TENT.
@@
We also test our method on CIFAR-10-C using another model ResNet50 (He et al. 2016). The results are shown in Table 5, which demonstrates that our methods can keep leading performance with various models compared to other methods w.r.t. average accuracy. Specifically, TTC is 2.3% better than TENT on CIFAR-10-C with Resnet50 and using gradient accumulation can get a 4.2% improvement.

### v6

@@ Abstract
Domain shift is a common problem in the realistic world, where training data and test data follow different data distributions. To deal with this problem, fully test-time adaptation (TTA) leverages the unlabeled data encountered during test time to adapt the model. In particular, entropy-based TTA (EBTTA) methods, which minimize the prediction’s entropy on test samples, have shown great success. In this paper, we introduce a new clustering perspective on the EBTTA. It is an iterative algorithm: 1) in the assignment step, the forward process of the EBTTA models is the assignment of labels for these test samples, and 2) in the updating step, the backward process is the update of the model via the assigned samples. This new perspective allows us to explore how entropy minimization influences test-time adaptation. Accordingly, this observation can guide us to put forward the improvement of EBTTA. We propose to improve EBTTA from the assignment step and the updating step, where robust label assignment, similarity-preserving constraint, sample selection, and gradient accumulation are proposed to explicitly utilize more information. Experimental results demonstrate that our method can achieve consistent improvements on various datasets. Code is provided in the supplementary material.
@@ Introduction
Despite the success of EBTTA, these previous studies still leave an open question: why or in which case the entropy minimization (Wang et al. 2021) can improve the TTA. Simply reducing the entropy may not always succeed. For example, the performance of EBTTA would still vary greatly across different situations. Zhao et al. (Zhao et al. 2023) had shown that TENT holds empirical sensitivity and batch dependency through extensive experiments. In Table 1, it even performs much worse than the source model when the batch size is 1. Hence, it is necessary to gain a more thorough understanding of EBTTA, and then come up with a better solution.
In this paper, we provide a new perspective to better understand EBTTA through a clustering view(Guo et al. 2017). The entropy-based methods and clustering both belong to unsupervised learning, where entropy minimization tries to reduce uncertainty about which category the sample belongs to, and clustering aims to map data into different clusters. Observing that, we show that entropy minimization and clustering have a lot in common, which makes it possible to use a clustering view to revisit the EBTTA. Based on the clustering view, the key components of entropy-based methods are not only at minimizing the entropy loss, but also at implicitly using assigned multiple samples to update the model. More specifically, we can view the forward process of the deep model in entropy minimization as the assignment step in clustering. According to Lemma 1, the entropy loss would increase the probability of the class with the largest value, which is similar to assigning the sample to its nearest clustering center. And the backward process (loss back-propagation and gradient descent) as the center updating step in clustering (see Section Interpretation for Entropy-Based Methods for detailed discussions). The entropy-based methods implicitly use the assigned multiple samples to update the model. This is the key to the success of the entropy-based methods.
Based on the interpretation, we can set the stage for improving the EBTTA. According to the clustering view, we improve the EBTTA from the assignment step and the updating step. In the assignment step, according to the observation in Lemma 1, the accurate assignment is very important for EBTTA. In this paper, we propose robust label assignment and similarity-preserving constraint to get a better initial assignment. The data augmentation is first used to obtain a robust initial probability for each test sample. Similarity-preserving constraint can promote that similar samples in feature space are assigned in the same cluster. In the updating step, the key is how to efficiently and stably use the assigned multiple samples to update the model. In this paper, we put forward to take advantage of the reliability of samples and the information from more samples, where sample selection and gradient accumulation are proposed to further improve the EBTTA. In summary, we propose to interpret the EBTTA from the view of clustering, which can guide us to develop better approaches for EBTTA.
@@
  • We revisit EBTTA from a clustering view: the forward process can be interpreted as the label assignment while the backward process of EBTTA can be interpreted as the updating of centers in clustering methods.
  • The clustering view can guide us to put forward the improvement of EBTTA, where the robust label assignment, similarity-preserving constraint, sample selection, and gradient accumulation are proposed to improve EBTTA.
  • We evaluate our method on various benchmark datasets, which demonstrates that our method can gain consistent improvements over existing EBTTA.
@@
Fully test-time adaptation (TTA) aims to solve the problem of domain shift, where the training (source) domain and the test (target) domain follow different distributions. In order to fulfill this target, TTA online adapts the model using test samples, and a large amount of TTA methods have been proposed. For example, T3A (Iwasawa and Matsuo 2021) and Wang et al. (Wang et al. 2023) proposed to compute prototypes for each class and online updates the prototypes during adaptations; TeSLA (Tomar et al. 2023), RoTTA (Yuan, Xie, and Li 2023) and RMT (Döbler, Marsden, and Yang 2023) enforced the consistency between the teacher and student model; TIPI (Nguyen et al. 2023) and MEMO (Zhang, Levine, and Finn 2022) proposed to learn a model invariant of transformations; SwapPrompt (Ma et al. 2024) explored the application of TTA on vision-language model.
The entropy-based test-time adaptation is a simple yet effective practice in the test-time adaptation, which minimizes the prediction’s entropy of test samples to adapt the model at test time. The most representative work of EBTTA is TENT (Wang et al. 2021), which conducted entropy minimizing via stochastic gradient descent and only modulated the BN layers. Then, several methods have been proposed to improve TENT. For example, ETA (Niu et al. 2022) and SAR (Niu et al. 2023) also minimized prediction’s entropy but only selects the reliable samples as inputs. SHOT (Liang, Hu, and Feng 2020) augmented TENT by adding a divergence loss. AdaContrast (Chen et al. 2022) uses nearest neighbor information to generate more correct pseudo labels. CoTTA (Wang et al. 2022) used TENT to implement continual TTA, where the target domain changes sequentially. Despite the success of EBTTA, several works have found that the performance of EBTTA may sharply drop in some cases, e.g., when adapting on high-entropy samples (Niu et al. 2022) or when batch size is small (Wang et al. 2021; Nguyen et al. 2023). To explain this, we propose to view EBTTA as a clustering method and further make improvements on EBTTA based on the clustering view.
@@
Deep Spectral Clustering
@@
The classical spectral clustering (Ng, Jordan, and Weiss 2001) stated above needs the calculation for eigenvectors thus lacks scalability. Deep spectral clustering (Shaham et al. 2018; Yang et al. 2019) is proposed to tackle this problem by learning the spectral embedding with a deep model using the locality-preserving constraint (Huang et al. 2014) as loss function. This approximation to classical spectral clustering can get better scalability and generalization ability with the cost of accuracy (Shaham et al. 2018).
@@
Mini-Batch K-Means: We give a brief description of mini-batch k-means (Bottou and Bengio 1994; Béjar Alonso 2013), which is used to interpret EBTTA methods. In the mini-batch k-means, the centers are first initialized, and then the A (assignment) and U (updating) steps are conducted alternately .
@@
U-step: With the probability $P(x)$ for each sample $x$, EBTTA uses the entropy loss to update the model. This process is similar to the clustering methods that update the centers. In the clustering methods, the $i$-th center is updated by averaging features of the samples that were assigned to the $i$-th center in the A-step. We first give a lemma, which shows that the entropy loss also aims to further increase the samples’ probability belonging to the $i$-th class, where these samples have the largest probability of the $i$-th class in the forward process. The backward process of EBTTA is very similar to the U-step of clustering, that is both methods use the assigned samples to update the model/centers.
@@
For sample $x$, the entropy loss $H(x)$ would increase the probability of the class with the largest value, and decrease the sum of probabilities of the other classes .
@@
Table 1: Accuracy(%) of TENT collapses when batch size is 1.
Remark 1: In the A-step, the optimal/minimum of the entropy loss does not mean a good result.
@@
As shown in Table  1, TENT collapses when batch size is 1 and is worse than SOURCE method even though the entropy loss arrives minimum. Lemma 1 shows that if we get only one sample for each iteration (batch size = 1), then the model only tends to make the largest output probability predicted by source model larger until the output probability becomes one-hot as Figure 1 shows. More severely, the performance is even worse (will be discussed in Remark 2). Thus, what is truly important to the A-step?
@@
Q1: What important information does the A-step provide during adaptation?
@@
From the clustering view in the TTA method, the A-step is used to assign the mini-batch samples to the $K$ classes. As we had discussed above, the entropy loss tries to further increase the largest probability of each test sample. Thus the accurate assignment is very important, and we can achieve excellent performance with good assignment, otherwise minimizing the entropy loss may hurt the performances. This is similar to clustering methods that are also sensitive to the initial assignment.
@@
Several TTA works like AdaContrast (Chen et al. 2022) and TAST (Jang, Chung, and Chung 2023) use nearest neighbor information to get more accurate labels. These works are under the assumption that a sample and its nearest neighbor in the feature space are likely to share the same label under the domain shift. The effectiveness of these works can be easily explained from the view of spectral clustering which also considers nearest neighbor information: using spectral clustering can make the clusters robust to complex scenarios (Liu and Han 2018). For example, if data points in a cluster after domain shift follow a non-Gaussian distribution, then it would be a good choice to use spectral clustering to generate clusters.
In summary, from the clustering view, the truly important thing of the A-step is assigning the samples to the cluster centers as accurately as possible.
Remark 2: In the U-step, the good and stable cluster centers are updated via the assigned multiple samples but not single sample.
When batch size is large than 1, the gradients of multiple samples from target domain will be averaged to update a shared model. Hence, the largest probability of the data point may be changed due to the averaged gradients. This is similar to the clustering method. For example, $x_{1}$ is assigned to the positive class in the current iteration. After the update of cluster centers, $x_{1}$ may possibly be assigned to the negative class in the next iteration. This implicit use of multiple samples is an important step for EBTTA.
Q2: How to efficiently and stably use the assigned multiple samples to update the model?
@@
Several methods, e.g., ETA (Niu et al. 2022) and SAR (Niu et al. 2023), had proposed to select only the low-entropy samples to efficiently update the model, and the high-entropy samples were removed. The efficiency of this technique has been verified by experiments but a detailed analysis is missing. From the clustering view, we can easily give an explanation: outliers should be suppressed.
@@
The number of the multiple samples is the batch size. The EBTTA methods are also sensitive to the batch size (Wang et al. 2021; Nguyen et al. 2023). This can also be easily answered using the clustering view. As the deep model is updated via the averaged assigned samples’ gradients, more samples can provide more stable gradients. That is, when the batch size is small, the variance will be large. This is similar to the clustering methods, where the centers are updated by averaging features of assigned samples. When the number of samples is small, the calculated centers also will have large variance and become unstable.
From the view of clustering and the existing methods, using sample selection and larger number of samples become common practices.
@@
# Similarity-Preserving Constraint
@@
Test-Time Clustering (TTC)
Based on the clustering view and the above discussions, we propose to improve TENT from these two aspects accordingly. In the A-step, we suggest to obtain a more robust initial label assignment in order to get high-quality initial probabilities. We also propose to add the similarity-preserving constraint to approximate spectral clustering, which improves the assignment quality. We use these two modules to obtain more accurate assignments. In the U-step, we propose to select test samples to alleviate the negative effects of outliers Finally, we recommend to use gradient accumulation to overcome the problem of small batch size.
Robust Label Assignment (RLA)
As discussed in Remark 1, we can see that label assignment in the A-step is very important for the EBTTA, which also provides a good initial probability for each test sample.
$\bar{f}(x)=\frac{f(x)+f(aug(x))}{2}$ (7)
$aug(x)=horizontal\_flip(x).$ (8)
This is because a weakly-augmented unlabeled sample can preserve more information about class label than the strongly augmented functions (Yang et al. 2022).
Now the entropy loss for sample $x$ can be calculated as:
Similarity-Preserving Constraint (SPC)
Spectral clustering is an improvement of k-means, which considers the similarity between samples in the same class. It uses the spectral embedding as data points for cluster partitioning. We add the similarity-preserving constraint to generate spectral embedding (Huang et al. 2014; Yang et al. 2019; Shaham et al. 2018). Specifically, the constraint loss for a single sample $x$ can be formulated as:
Sample Selection (SS)
$H_{m}=\frac{1}{N}\sum_{x\in X}H(x).$ (11)
The selecting function $\mathcal{S}(x)$ can be formulated as:
Gradient Accumulation (GA)
The pseudo-code of our method is shown in Algorithm 1.
@@
The comparison results are shown in Table 2, Table 3 and Table 4. The average (avg) accuracy is calculated on all corruption types. The results show that our method can achieve consistent improvements for EBTTA methods w.r.t. average accuracy on all three datasets. More detailed analysis is as follows.
CIFAR-10-C. The results are shown in Table 2. We can see that TTC achieves the best average accuracy 82.5%, which is 1.0% superior to the second-best method TENT. Besides, TTC performs better than TENT on most corruption types. Specifically, TTC is 2.7% superior to TENT w.r.t. impulse noise, which is a non-Gaussian transformation. Since TENT is the most relevant work, the comparison results indicate that the proposed improvements are effective.
CIFAR-100-C. As shown in Table 3, TTC is 1.8% better than TENT. It’s noting that NORM has achieved a good average accuracy, and our method method has a very close result.
ImageNet-C. The results of Table 4 show that TTC achieves the best average accuracy 45.0%, and the accuracy of the second-best model TENT is 41.8%. Note that previous entropy-based methods achieve excellent performance, even so, our method performs better than them.
@@
We conduct experiments to analyze the improvements caused by different components of our method individually. Specifically, we first use TENT as the base model. We apply different components individually to TENT and the average accuracies for datasets are shown in Table 7. We shall see that robust label assignment (RLA) can bring about improvements for TENT on CIFAR-10-C and CIAFR-100-C. Similarity-preserving constraint (SPC) can improve the performance on CIFAR-100-C and ImageNet-C. Sample selection (SS) can bring about improvements for TENT on both CIFAR-10-C and ImageNet-C. Gradient accumulation (GA) can make significant improvements on CIFAR-100-C. Please note that the batch size in Table 7 is 100, which is not too small. Thus it may be the reason why TENT+GA has similar result to TENT on CIFAR-10-C and ImageNet-C dataset. Detailed analysis of GA for different batch sizes can be found in Section Analysis of Different Batch Sizes, where GA can boost the performance significantly when the batch size is small.
@@
We conduct experiments to explore the effect of different augmentation functions $aug(x)$. The augmentation functions are horizontal flip (HorFlip), vertical flip (VerFlip), random rotation (RandRot), random resized crop (RRCrop), and Gaussian noise (GausNoise). From the average accuracies for datasets in Table 5, we shall see that horizontal flip makes the performance better than other augmentation functions applied to TTC.
@@
We further conduct an ablation study to explore the effect of the batch size, and the average accuracies on datasets are shown in Figure 2. We shall see that all methods perform better when batch size increases. Moreover, when applying gradient accumulation (GA) for small batch sizes, we can get a large performance improvement for TENT. For example, when the batch size is 2, TENT only gets 17.1% accuracy on CIFAR-10-C. However, when applying GA, we can get an improvement of 43.6% on TENT.
@@
We also test our method on CIFAR-10-C using another model ResNet50 (He et al. 2016). The results are shown in Table 6, which demonstrates that our methods can keep leading performance with various models compared to other methods w.r.t. average accuracy. Specifically, TTC is 2.9% better than TENT on CIFAR-10-C with Resnet50.